show that ⠈ ‘ (-1)^(n+1)à —5à ·(7n+2) is condintionally convergent
1
Expert's answer
2021-10-12T15:42:10-0400

Since, given term is unclear, we assume it as "\sum _{n=0}^{\infty \:}\frac{\left(-1\right)^{n+1}\cdot \:5}{7n+2}"

Solution:

"\sum _{n=0}^{\infty \:}\frac{\left(-1\right)^{n+1}\cdot \:5}{7n+2}"

"=5\cdot \sum _{n=0}^{\infty \:}\frac{\left(-1\right)^{n+1}}{7n+2}"

"=5\cdot \sum _{n=0}^{\infty \:}\frac{\left(-1\right)^n\left(-1\right)^1}{7n+2}"

"=5\left(-1\right)\cdot \sum _{n=0}^{\infty \:}\frac{\left(-1\right)^n}{7n+2}"

Apply alternating series test

(a) To check "\lim_{n\rightarrow \infty} b_n=0"

"\lim_{n\rightarrow \infty} \dfrac{1}{7n+2}=\dfrac1{\infty}=0"

(b) And to check "b_{n+1}\le b_n"

"b_n=\dfrac1{7n+2}, b_{n+1}=\dfrac1{7(n+1)+2}=\dfrac1{7n+9}"

Clearly, "\dfrac1{7n+9}\le \dfrac1{7n+2}"

Thus, "b_{n+1}\le b_n"

Hence, by using alternating series test

"\sum _{n=0}^{\infty \:}\frac{\left(-1\right)^{n+1}\cdot \:5}{7n+2}" "=5\left(-1\right)\mathrm{converges}"

"=\mathrm{converges}"